Please help me with this problem.
This spinner is spun 36 times. The spinner landed on A 6 times, on B 21 times, and on C 9 times. Compare the empirical probability that the spinner will land on B.
Please show step by step so I can better understand.

Answer by richard1234(7193)

The empirical probability is 21/(6+21+9) = 21/36 = 7/12.